OpenStudy (anonymous):

The population (in millions) of a certain country can be approximated by the function: P(x)=50*1.02^x where x is the number of years after 2000. Which of the following calculations will tell in what year the population can be expected to reach 100 million? a. ln(2/1.02)+2000 b. ln(2)/ln(1.02)+2000 c. ln(2/1.02) d. ln(2)/ln(1.02)

OpenStudy (anonymous):

\[100=50(1.02)^x\] \[2=(1.02)^x\] \[x=\frac{\ln(2)}{\ln(1.02)}\]

OpenStudy (anonymous):

i guess you have to add 2000 to get the actual year

OpenStudy (anonymous):

wow thank you so much, you are amazing!!

OpenStudy (anonymous):

so strange as it seems, go with B

OpenStudy (anonymous):

yw